For the a Poisson random variable x generating function G(t)= exp^lambda(t-1) show that the expected value E(x) and variance (x) are given by lambda and lambda respectively.

1
Expert's answer
2022-01-17T12:37:37-0500

"G(t)=e^{\lambda(t-1)}"

We want to verify that "var(x)=E(x)=\lambda"

Now,

"E(x)=G'(t=1)"

First,

"G'(t)=\lambda e^{\lambda(t-1)}"

Setting "t=1" in "G'(t)",

"G'(t=1)=\lambda e^{\lambda\times0}=\lambda\times1=\lambda"

Therefore, "E(x)=\lambda"

To find "var(x)", we first determine "E(x^2)-E(x)=G''(t)"

Now,

"G''(t)=\lambda^2e^{\lambda(t-1)}"

Setting "t=1" in "G''(t=1)" we get,

"G''(t=1)=\lambda^2e^{0}=\lambda^2\times1=\lambda^2"

We determine the value of "E(x^2)" as follows.

"G''(t)=E(x^2)-E(x)\implies E(x^2)=G''(t)+E(x)=\lambda^2+\lambda"

So,

"E(x^2)=\lambda^2+\lambda"

Thus,

"var(x)=E(x^2)-(E(X))^2\\
var(x)=(\lambda^2+\lambda)-\lambda^2=\lambda"

Therefore, "E(x)=var(x)=\lambda" as required.